# Transformation-free linear simplicial-simplicial regression via constrained iterative reweighted least squares

### v1

@@ 4.2 Scalability of the EM and CIRLS algorithms
To further examine and compare the scalability of the two algorithms, we estimated their scalability rate by fitting the following model $t=\alpha n^{\beta}$, or $\log(t)=\alpha+\beta n$ (goldsmith2007), and report the $\beta$ coefficient, where $t$ and $n$ denote the running time and sample size. Table 1 presents the empirically estimated scalability rates of both algorithms. Interestingly enough, the CIRLS possess a linear scalability rate, whereas the EM has a sub-linear rate, explaining why the speed-up factors decrease with increasing sample size.
Table 1: Scalability rates of the EM and CIRLS algorithms.
Algorithm D=3 D=5 D=7 D=10
$p=5$ EM 0.32 0.26 0.31 0.34
CIRLS 0.94 0.91 1.02 0.98
$p=10$ EM 0.22 0.46 0.53 0.62
CIRLS 1.00 0.95 0.90 0.92
@@ 5 Conclusions
We adopted the IRLS methodology used in generalised linear models to the case of the TFLR model for simplicial-simplicial regression. We exploited the fact that TFLR may be seen as a logistic regression model with simplicial constraints on the regression coefficients. The time improvements were significant, ranging from 2 up to 85 times. The estimated scalability rates though showed that the EM has sub-linear rate, whereas the CIRLS has a linear rate.

### v2

@@ 5.2 Scalability of the EM and CIRLS algorithms
To further examine and compare the scalability of the two algorithms, we estimated their scalability rate by fitting the following model $t=\alpha n^{\beta}$, or $\log(t)=\alpha+\beta n$ (goldsmith2007), and report the $\beta$ coefficient, where $t$ and $n$ denote the running time and sample size. Tables 2 and 3 present the empirically estimated scalability rates of both algorithms. Notably, all implementations exhibit linear or sublinear scalability rates, depending on the dimensionalities of the compositions.
Table 2: Estimated scalability rates of the EM and CIRLS algorithms when the responses and the predictor variables are independent.
Algorithm $D_{r}=3$ $D_{r}=5$ $D_{r}=7$ $D_{r}=10$
$p=5$ EM-1 0.90 0.87 0.89 0.89
EM-2 0.21 0.28 0.40 0.42
CIRLS 0.81 1.14 0.77 1.16
$p=10$ EM-1 0.79 0.72 0.76 0.73
EM-2 0.13 0.63 0.68 0.73
CIRLS 0.96 0.81 1.07 1.11
$p=15$ EM-2 0.66 0.48 0.36 0.30
CIRLS 0.82 0.70 0.60 0.54
$p=20$ EM-2 0.32 0.14 0.11 0.17
CIRLS 0.66 0.52 0.45 0.36
Table 3: Estimated scalability rates of the EM and CIRLS algorithms when the responses and the predictor variables are dependent.
Algorithm $D_{r}=3$ $D_{r}=5$ $D_{r}=7$ $D_{r}=10$
$p=5$ EM-1 0.92 0.89 0.76 0.77
EM-2 1.01 0.91 1.03 0.99
CIRLS 0.64 0.98 0.98 0.87
$p=10$ EM-1 0.82 0.77 0.69 0.63
EM-2 1.03 1.01 0.95 0.87
CIRLS 0.88 0.98 0.97 0.90
$p=15$ EM-2 0.38 0.52 0.48 0.55
CIRLS 0.77 0.77 0.62 0.47
$p=20$ EM-2 0.73 0.66 0.67 0.64
CIRLS 0.78 0.62 0.44 0.26
@@ 6 Conclusions
We adopted the IRLS methodology used in generalised linear models to the case of the TFLR model for simplicial-simplicial regression. For this we exploited the fact that TFLR may be seen as a logistic regression model with simplicial constraints on the regression coefficients. Regardless of the dimensiolnalities in both the responses and the predictors, this resulted in significant time improvements ranging from 1.3 up to more than 300 times. The estimated scalability rates were linear or sublinear for both algorithms.
The impressive speed-up factors of CIRLS though come with a cost. When the compositions are dependent, the CIRLS fails by a small amount ($\approx 10^{-4}$) to reach the minimum KLD value. Our suggested compromise is to perform the CIRLS and use the estimated regression coefficients as initial values in the EM algorithm. For bootstrap or permutation hypothesis testing though, the CIRLS might be a good alternative.